# Articulated Object Manipulation using Online Axis Estimation with SAM2-Based Tracking

### v1

@@ IV EXPERIMENT
IV-A Task Settings
Our method specifically focuses on door-opening and drawer-opening tasks. We evaluate the performance of these tasks in scenarios with large ranges of door opening and drawer extension.
In each task, the articulated objects, placed with limited random position and rotation, begins in a closed state, and the robotic arm is initially positioned randomly in front of it. The arm must then accomplish the designated manipulation goal — either opening the drawer or door to a specific degree or range.
The task settings are detailed as follows:
Figure 3: Success rate of more challenging tasks for opening door.
Figure 4: Success rate of more challenging tasks for opening drawer.
@@ IV-A Task Settings
Methods Open Door 8.6° Open Door 45° Open Drawer 15 cm Open Drawer 30cm
@@
DrQ-v2  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [5]. 1.8 2.5 0.8 0.8 1.9 1.0 1.4 0.5
LookCloser  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [5]. 1.5 1.25 0.8 0.8 0.8 0.0 0.0 0.0
RGBManip  [^2] Results reproduced using RGBManip under the settings outlined in Sec. IV-A. 75.0 82.0 47.0 47.0 56.0 64.0 46.0 45.0
Where2Act  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [5]. 8.0 7.0 1.8 2.0 5.9 7.5 1.1 0.6
Flowbot3D  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [5]. 19.5 20.4 6.8 6.4 27.3 25.8 16.9 11.3
UMPNet  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [5]. 27.1 28.1 11.0 10.9 16.6 18.8 4.4 5.6
GAPartNet  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [5]. 69.5 74.5 39.4 43.6 50.6 59.3 44.6 48.6
Ours 87.0 88.0 54.0 54.0 68.0 85.0 59.0 68.0
@@
IV-B Baselines and Quantitative Results
@@
Fine-grained results for more challenging drawer-opening tasks are provided in Table. III, with the corresponding line chart illustrating success rates at various drawer-opening distances shown in Fig. 4. Experimental results show that Ours consistently outperforms RGBManip with a significant enhancement in success rates.
@@ V Analysis
Furthermore, in the more demanding versions of these tasks, depicted in Figures 3 and 4, our method’s reliance on ongoing online axis estimation reveals enhanced robustness in handling large-amplitude movements of articulated objects. The experimental data consistently show that our method sustains higher success rates as the complexity of tasks escalates, whether it involves opening doors to wider angles or drawers to greater extents. The online axis estimation process is pivotal in this regard, enabling our system to dynamically adjust to real-time interaction feedback and ensuring precise control over articulated objects. The continuous axis refinement, aligned with the object’s changing state, is essential for the manipulation’s accuracy and efficiency, particularly in tasks with substantial state variations. The real-time feedback loop is crucial, enabling the robot to make swift and accurate operational adjustments that accommodate the subtleties of the interaction, thereby achieving a higher success rate in task completion. This robust performance underscores the strength of online axis estimation in providing reliable and responsive control in sophisticated robotic manipulation scenarios.
@@ VI CONCLUSION
In this research, we introduced a novel approach to articulated object manipulation that employs online axis estimation integrated with SAM2-based tracking. Our method utilizes segmented 3D point cloud data and motion axis estimation to enhance the precision and control of robotic actions. The experimental results suggest that our method shows promising improvements in handling large manipulation movements, indicating the potential utility of axis-based guidance in robotic manipulation tasks.
The online axis estimation technique has demonstrated beneficial in dynamically adjusting to the real-time interaction feedback, thereby enabling the robot to maintain better control over articulated objects. While our method has shown enhancements over traditional approaches, particularly in complex manipulation scenarios, it is important to acknowledge that there is still room for further refinement and optimization.

### v2

@@ IV EXPERIMENT
Methods Modality Open Door 8.6° Open Door 45° Open Drawer 15 cm Open Drawer 30cm
@@
DrQ-v2  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [12]. RGB 1.8 2.5 0.8 0.8 1.9 1.0 1.4 0.5
LookCloser  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [12]. RGB 1.5 1.25 0.8 0.8 0.8 0.0 0.0 0.0
RGBManip  [^2] Results reproduced using RGBManip under the settings outlined in Sec. IV-A. RGB 75.0 82.0 47.0 47.0 56.0 64.0 46.0 45.0
[…]
Flowbot3D  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [12]. PCD 19.5 20.4 6.8 6.4 27.3 25.8 16.9 11.3
UMPNet  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [12]. PCD 27.1 28.1 11.0 10.9 16.6 18.8 4.4 5.6
GAPartNet  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [12]. PCD 69.5 74.5 39.4 43.6 50.6 59.3 44.6 48.6
[…]
@@
IV-A Task Settings
In this study, we specifically focus on 1-DoF doors and drawers for experimental validation. We evaluate the performance in scenarios with large ranges of door-opening and drawer-extension.
In each task, the articulated objects, placed with limited random position and rotation, begins in a closed state, and the robotic arm is initially positioned randomly in front of it. The arm must accomplish the designated manipulation goal, either opening the drawer or door to a specific degree or range.
The task settings are detailed as follows:
We select $r=0.05$ and $\epsilon=100$ as the hyperparameters for point cloud augmentation.
IV-B Baselines
@@ IV-C Quantitative Results
  • Fine-grained results for more challenging drawer-opening tasks are provided in Table. III, with the corresponding line chart illustrating success rates at various drawer-opening distances shown in Fig. 5.
Experimental results show that Ours consistently outperforms RGBManip with a significant enhancement in success rates, showing robustness across various types of cabinets.
IV-D Real-World Deployment
Figure 3: Visualization of axis estimation for real-world “Open Drawer” and “Open Door” tasks. The initial moment and the three manipulation moments are shown, with visualization of the RGB tracking obtained from SAM2 (background), the reconstructed point cloud of the target object (bottom-left corner), the OBBs (green dashed-line boxes at $t_{0}$) and the axis (red arrows) estimated with our method.
To validate the effectiveness of our method in real-world deployment, we demonstrate the complete axis estimation process for the door-opening and drawer-opening tasks.
We employ the D415 depth camera. The hyperparameters are set to $r=1.3$ and $\epsilon=1$, due to the sparsity of the point cloud reconstructed from the real-world depth map.
Empirical results are shown in Fig. 3. Despite the presence of noise in real-world point clouds, our method remains robust. This is attributes to our point cloud augmentation method, which enables robust computation of OBBs, leading to accurate and reliable axis estimation in noisy scenarios.
@@ V Analysis
Figure 4: Success rate of more challenging tasks for opening door.
Figure 5: Success rate of more challenging tasks for opening drawer.
@@
Furthermore, in the more demanding versions of these tasks, depicted in Figures 4 and 5, our method’s reliance on ongoing online axis estimation reveals enhanced robustness in handling large-amplitude movements of articulated objects. The experimental data consistently show that our method sustains higher success rates as the complexity of tasks escalates, whether it involves opening doors to wider angles or drawers to greater extents. The online axis estimation process is pivotal in this regard, enabling our system to dynamically adjust to real-time interaction feedback and ensuring precise control over articulated objects. The continuous axis refinement, aligned with the object’s changing state, is essential for the manipulation’s accuracy and efficiency, particularly in tasks with substantial state variations. The real-time feedback loop is crucial, enabling the robot to make swift and accurate operational adjustments that accommodate the subtleties of the interaction, thereby achieving a higher success rate in task completion. This robust performance underscores the strength of online axis estimation in providing reliable and responsive control in sophisticated robotic manipulation scenarios.
@@ VI CONCLUSION
In this research, we introduced a novel approach to articulated object manipulation that employs online axis estimation integrated with SAM2-based tracking. We utilizes segmented 3D point cloud data and motion axis estimation to enhance the precision and control of robotic actions. Experimental results suggest that our method shows promising improvements in handling large manipulation movements, indicating the potential utility of axis-based guidance in robotic manipulation tasks.
The online axis estimation technique has demonstrated beneficial in dynamically adjusting to the real-time interaction feedback, thereby enabling the robot to maintain better control over articulated objects. While our method has shown enhancements over traditional approaches, there is still room for further refinement and optimization.